A pocket contains nickels and dimes. The value of the coins is 90 cents. The number of nickels is 2 less than three times the number of dimes. Find the number of Nickels in the pocket
Answer by TimothyLamb(4379)

x = number of nickels
y = number of dimes
---
5x + 10y = 90
x = 3y - 2
---
put the system of linear equations into standard form
---
5x + 10y = 90
x - 3y = -2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 10
y = 4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---